OpenStudy (anonymous):

-5(n+3)+4(1+5n)=4 whats n

jimthompson5910 (jim_thompson5910):

-5(n+3)+4(1+5n)=4 -5n-15+4+20n=4 15n-11=4 15n = 4 + 11 15n = 15 n = 15/15 n = 1